# Quantum Algorithm For Solving Nonlinear Algebraic Equations

### v1

@@ Abstract
Nonlinear equations are challenging to solve due to their inherently nonlinear nature. As analytical solutions typically do not exist, numerical methods have been developed to tackle their solutions. In this article, we give a quantum algorithm for solving a system of nonlinear algebraic equations, in which each equation is a multivariate polynomial of known coefficients. Building upon the classical Newton method and some recent works on quantum algorithm plus block encoding from the quantum singular value transformation, we show how to invert the Jacobian matrix to execute Newton’s iterative method for solving nonlinear equations, where each contributing equation is a homogeneous polynomial of an even degree. A detailed analysis are then carried out to reveal that our method achieves polylogarithmic time in relative to the number of variables. Furthermore, the number of required qubits is logarithmic in the number of variables. In particular, we also show that our method can be modified with little effort to deal with polynomial of various types, thus implying the generality of our approach. Some examples coming from physical context, such as Gross-Pitaevski equation and Lotka-Volterra equations, involving nonlinear partial differential equations are provided to motivate the potential application, with a description on how to extend our algorithm with even less effort in such a scenario. Our work thus marks a further important step towards quantum advantage in nonlinear science, enabled by the framework of quantum singular value transformation.
@@ IV Block Encoding the Jacobian J and its inversion
$M_{D,j}=\begin{pmatrix}M_{D,j}^{1}&\cdot&\cdot&\cdot\\ \cdot&M_{D,j}^{2}&\cdot&\cdot\\ \cdot&\cdot&\dots&\cdot\\ \cdot&\cdot&\cdot&M_{D,j}^{n}\\ \end{pmatrix}$
@@
We remark that the complexity $\mathcal{O}(1)$ appeared above does not include the complexity of preparing the block encoding of $P/(ps)$, and the value of $\gamma$, while being bounded, changes over different iteration steps. Furthermore, the factor ${\gamma^{2p-1}}/({\sqrt{n}ps})$ in the above formula is seemingly unfavorable. Fortunately, we can remove the factor $ps$ using the technique called preamplification gilyen2019quantum . Such a method requires the usage of the unitary $U$ that encodes the operator $\hat{O}_{J}$ roughly $\mathcal{O}(\sqrt{ps})$ times. Therefore, the total time required to prepare the $\epsilon$-approximated block encoding of $\gamma^{2p-1}J/\sqrt{n}$ is
$\mathcal{O}\Big{(}\sqrt{ps}\ (\mathcal{T}+p(\log(n)+\log^{2.5}(\frac{1}{\epsilon})))\Big{)}.$ (19)
@@
$\mathcal{O}\Big{(}\sqrt{ps}(\mathcal{T}+p(\log(n)+\log^{2.5}(\frac{1}{\epsilon})))\cdot\frac{1}{\sigma}\cdot polylog(\frac{1}{\sigma\epsilon})\Big{)},$
where $s$ is the sparsity as defined, and $\sigma$ is some constant that is at least greater than the smallest singular value of $\gamma^{2p-1}J/\sqrt{n}$
@@ V Preparing Right-Hand Side State
From this we perform the following steps:
@@
Using the preamplification technique, one can remove the factor $|x|^{2}$ to obtain the block encoding of $F(x)F(x)^{T}$ with a further time complexity $\mathcal{O}(1/|x|^{2})$. The remaining question now is, in order to remove such a factor $|x|^{2}$, we need to know the value. This is exactly the norm of the solution, and it can be estimated using Lemma 9. To show this, we simply make use of the following observation: at a given iteration step, we are presented with the block encoding of $xx^{T}$. This operator has a special property that $|x|^{2}$ is its largest eigenvalue. Therefore, Lemma 9 is sufficient to estimate such factor $|x|^{2}$.
@@ VI Quantum Newton Method
Before heading to the final stage, we emphasize an important point that accounts for the reason why we keep the factor $\gamma^{2p-1}/\sqrt{n}$ throughout the work. The above formula contains the term $J^{-1}F(x)x$ (where $x$ refers to any input vector). The inversion of $\gamma^{2p-1}J/\sqrt{n}$ is $\sqrt{n}J^{-1}/\gamma^{2p-1}$, when it multiplies with the block encoding of $\gamma^{2p-1}F(x)x^{T}/\sqrt{n}$, the unwanted factors will cancel out. Therefore, we do not need to use the preamplification technique to remove the factor before executing the algorithm, which results in a very efficient procedure.
@@
Therefore, once we obtain the block encoding of the above operator, we produce the block encoding of $x_{k+1}x_{k+1}^{T}$ with a further $\mathcal{O}(\sigma)$ time complexity.
@@
$\mathcal{O}\Big{(}\sqrt{ps}\big{(}\log(n)+\frac{1}{\sigma_{k}}{\rm polylog}(\frac{1}{\sigma_{k}\epsilon})\big{)}\big{(}\mathcal{T}_{k}+\log(n)+\log^{2.5}(1/\epsilon)\big{)}\Big{)}.$
@@
$\mathcal{O}\Big{(}\sqrt{ps}^{T+1}\big{(}\log(n)+\frac{1}{\sigma_{min}}{\rm polylog}(\frac{1}{\sigma_{min}\epsilon})\big{)}^{T+1}\big{(}\log(n)+\log^{2.5}({1}/{\epsilon})\big{)}\Big{)},$
@@ VII.1 Classical Algorithm and Quantum Advantage
$D_{i}(x)\cdot x=\sum_{\alpha=1}^{K}\sum_{m=1}^{p}\Big{(}\prod_{n=1,n\neq m}^{p}x^{T}A_{n}^{\alpha,i}x\Big{)}A_{m}^{\alpha,i}\cdot x,$
which takes a further complexity $\mathcal{O}(Kp^{2}Ns)$. Finally, solving the linear equation Eqn. (4) takes a further $\mathcal{O}(N^{3})$ time. Therefore, summing up the steps, the total time required for a total of $T$ iteration steps is:
$\mathcal{O}\Big{(}T(N^{3}+Kp^{2}Ns+N^{p+1}s)\Big{)}.$
For large $p$, e.g., $p\geq 3$, the task is practically intractable on a classical computer. The above running time indicates that the speedup of the quantum algorithm is exponential with respect to the number of variables and the degree level $p$, which is substantial. Furthermore, the quantum running time depends on the behavior of the Jacobian as well, i.e., its spectrum over iteration steps. It is generally a non-trivial problem to determine the behavior of the landscape of the Jacobian’s spectrum over some domain.
As we mentioned, it is well-known that Newton’s method typically works well given that we initially randomize a vector that is within the vicinity of the real solution. Therefore, it is reasonable that Newton’s method (both classical and the quantum versions) will succeed after a few iterations, e.g., when $T$ is small. In this case, a huge quantum speedup is obtained. Now, we move on to discuss a way to achieve better initialization from a quantum algorithmic perspective.
@@ VII.2 Quantum Approach for Initialization
Aside from the particular algebraic form that we impose, the behavior of functions of interest is quite critical. We remind that the Newton method is usually effective when the initially randomized solution is closed to the real one, as the method would succeed after after a few iterations. Therefore, it is quite important to sort out a way to prepare such a good initialization. A general strategy to prepare such a thing is not known to us. Here we aim to propose a way to accelerate such a task, inspired by the way that each of the function $f_{i}$ can be computed quickly using quantum computer.
@@ VIII Motivating Examples And Potential Applications/Extensions
Nonlinear Optical Process
@@
where $\Delta t,\Delta x$ and $V_{j}^{n}$ are time step, grid spacing, and potential at the grid point $(x_{j},t_{n})$. The above equation at different grid points represents a coupled nonlinear algebraic equation, which could be solved using the classical Newton method to obtain a vector solution that contains all $\{\psi_{j}^{n}\}$ at different grid points. However, we point out two issues. First, our method is only applied to a nonlinear system where each equation is a homogeneous polynomial of even degree. Second, the above equation (Eqn. 50) actually contains both linear and nonlinear terms. In order to integrate our method, an adjustment is hence required. Apparently in Section VII.3 we showed how to deal with this scenario, but here we point out that it can be alternatively done in a simpler manner.
@@
$g=\begin{pmatrix}g_{1}=a_{1}x+b_{1}y+c_{1}z\\ g_{2}=a_{2}x+b_{2}y+c_{2}z\\ g_{3}=a_{3}x+b_{3}y+c_{3}z\end{pmatrix}$ (63)
@@
Now, we observe that the nonlinear equation in Ref. 50 contains both linear and nonlinear terms. Therefore, if we separate the function on the left-hand side of Eqn. (50) into linear and nonlinear parts, then the Jacobian of such equations, which basically contains gradients of corresponding functions, is the sum of the Jacobians from the constituting functions. The Jacobian of the linear constituent function is easily computed based on the above example. The Jacobian of the nonlinear constituent is computed using the algorithm outlined in our discussion above.
@@
Lotka-Volterra Equations for Modeling
@@ IX Conclusion
In this article, we have developed a quantum algorithm for solving nonlinear algebraic equations of arbitrary polynomial type. By employing well-known quantum tools, such as the unitary block encoding method, an iterative procedure building upon classical Newton’s method is constructed to drive the initially randomized guess to the close-to-real solution, which is expected to succeed after only a few iteration steps. Additionally, we have provided two examples from physical and biophysics contexts to suggest the potential application of our method, as well as suggesting another way to extend our method to deal with a more specific setting.
@@ X Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under Award Number DE-SC-0012704. We also acknowledge the support by a Seed Grant from Stony Brook University’s Office of the Vice President for Research and by the Center for Distributed Quantum Processing.
@@ Appendix C Proof of Lemma 7
According to the definition 1, the unitary $U_{m}U_{p}$ is the block encoding of the matrix:
$\frac{x_{0}^{2p-1}J}{\sqrt{n}ps}.$
One can reasonably expect that if the value of $x_{0}$ is very small, then it will incur a substantial running time. A better way to overcome this issue is to change the state |$\ket{\bf 0}_{A}$ in the above construction to some states that are close to any temporal solution $x_{k}$ during the iteration. As we have mentioned, suppose the initially guessed solution $x_{0}$ is within the vicinity of the actual solution, then the overlaps between any pairs of $x_{0}$ and temporal solution $x_{k}$ (at $k$-th iteration step) is supposed to be sufficiently high, e.g, lower bounded. Therefore, in the above procedure, one simply needs to let unitary $U$ acts on the register $A$ to transform the state |$\ket{\bf 0}_{A}$ to the state |$\ket{x_{0}}$ (or $x_{0}$ as we denoted before) and proceed accordingly. Then the factor $\gamma$ in Lemma 7 would be the overlap $|x_{0}^{T}x_{k}|$, which could be estimated using amplitude estimation [29, 30]. This step only incur a constant cost overhead and hence, don’t increase the overall running time.

### v2

@@ Abstract
Nonlinear equations are challenging to solve due to their inherently nonlinear nature. As analytical solutions typically do not exist, numerical methods have been developed to tackle their solutions. In this article, we give a quantum algorithm for solving a system of nonlinear algebraic equations, in which each equation is a multivariate polynomial of known coefficients. Building upon the classical Newton method and some recent works on quantum algorithm plus block encoding from the quantum singular value transformation, we show how to invert the Jacobian matrix to execute Newton’s iterative method for solving nonlinear equations, where each contributing equation is a homogeneous polynomial of an even degree. A detailed analysis are then carried out to reveal that our method achieves polylogarithmic time in relative to the number of variables. Furthermore, the number of required qubits is logarithmic in the number of variables. In particular, we also show that our method can be modified with little effort to deal with polynomial of various types, thus implying the generality of our approach. Some examples coming from physics and algebraic geometry, such as Gross-Pitaevski equation, Lotka-Volterra equations, and intersection of algebraic varieties, involving nonlinear partial differential equations are provided to motivate the potential application, with a description on how to extend our algorithm with even less effort in such a scenario. Our work thus marks a further important step towards quantum advantage in nonlinear science, enabled by the framework of quantum singular value transformation.
@@ I Introduction
To highlight our contribution, we emphasize that the number of qubits required, and the running time of our method is logarithmic and polylogarithmic, respectively, in relative to the number of variables in the algebraic equations. Further, our method could be applied to polynomial of any orders. Altogether, it implies the major improvement over previous attempts xue2021quantum ; xue2022quantum ; qian2019quantum , which is enabled by the recent striking quantum singular value transformation framework.
@@ IV Block Encoding the Jacobian J and its inversion
$M_{D,j}=\begin{pmatrix}M_{D,j}^{1}&\cdot&\cdot&\cdot\\ \cdot&M_{D,j}^{2}&\cdot&\cdot\\ \cdot&\cdot&\dots&\cdot\\ \cdot&\cdot&\cdot&M_{D,j}^{n}\\ \end{pmatrix}.$
@@
We remark that the complexity $\mathcal{O}(1)$ appeared above does not include the complexity of preparing the block encoding of $P/(ps)$, and the value of $\gamma$, while being bounded, changes over different iteration steps. Furthermore, the factor ${\gamma^{2p-1}}/({\sqrt{n}ps})$ in the above formula is seemingly unfavorable. Fortunately, we can remove the factor $ps$ using the technique called amplification gilyen2019quantum (Theorem 30 of gilyen2019quantum ). Such a method requires the usage of the unitary $U$ that encodes the operator $\hat{O}_{J}$ $\mathcal{O}(ps\log(ps/\epsilon))$ times. Therefore, the total time required to prepare the $\epsilon$-approximated block encoding of $\gamma^{2p-1}J/\sqrt{n}$ is
$\mathcal{O}\Big{(}ps\log(\frac{ps}{\epsilon})\ (\mathcal{T}+p(\log(n)+\log^{2.5}(\frac{1}{\epsilon})))\Big{)}.$ (19)
@@
$\mathcal{O}\left(ps\log(\frac{ps}{\epsilon})\Big{(}\mathcal{T}+p\big{(}\log(n)+\log^{2.5}(\frac{1}{\epsilon})\big{)}\Big{)}\cdot\frac{1}{\sigma}\cdot polylog(\frac{1}{\sigma\epsilon})\right),$
where $s$ is the sparsity as defined, and $\sigma$ is some constant that is at least greater than the smallest singular value of $\gamma^{2p-1}J/\sqrt{n}$.
@@ V Preparing Right-Hand Side State
From this, we perform the following steps:
@@
Using the amplification technique (essentially the uniform singular value transformation, Theorem 30 of gilyen2019quantum ), one can remove the factor $|x|^{2}$ to obtain an $\epsilon$-approximation block encoding of $F(x)F(x)^{T}$ with a further time complexity $\mathcal{O}(\log(|x|)/(|x|^{2}\epsilon))$. The remaining question now is, in order to remove such a factor $|x|^{2}$, we need to know the value. This is exactly the norm of the solution, and it can be estimated using Lemma 9. To show this, we simply make use of the following observation: at a given iteration step, we are presented with the block encoding of $xx^{T}$. This operator has a special property that $|x|^{2}$ is its largest eigenvalue. Therefore, Lemma 9 is sufficient to estimate such factor $|x|^{2}$.
@@ VI Quantum Newton Method
Before heading to the final stage, we emphasize an important point that accounts for the reason why we keep the factor $\gamma^{2p-1}/\sqrt{n}$ throughout the work. The above formula contains the term $J^{-1}F(x)x$ (where $x$ refers to any input vector). The inversion of $\gamma^{2p-1}J/\sqrt{n}$ is $\sqrt{n}J^{-1}/\gamma^{2p-1}$, when it multiplies with the block encoding of $\gamma^{2p-1}F(x)x^{T}/\sqrt{n}$, the unwanted factors will cancel out. Therefore, we do not need to use the amplification technique to remove the factor before executing the algorithm, which results in a very efficient procedure.
@@
Therefore, once we obtain the block encoding of the above operator, we produce the block encoding of $x_{k+1}x_{k+1}^{T}$ with a further $\mathcal{O}(\sigma_{k}\log(\sigma_{k}))$ time complexity gilyen2019quantum .
@@
$\mathcal{O}\Big{(}ps\log(\frac{ps}{\epsilon})\big{(}\log(n)+\frac{1}{\sigma_{k}}{\rm polylog}(\frac{1}{\sigma_{k}\epsilon})\big{)}\big{(}\mathcal{T}_{k}+\log(n)+\log^{2.5}(\frac{1}{\epsilon})\big{)}\Big{)}.$
@@
$\mathcal{O}\Big{(}(ps\log(\frac{ps}{\epsilon})))^{T+1}\big{(}\log(n)+\frac{1}{\sigma_{min}}{\rm polylog}(\frac{1}{\sigma_{min}\epsilon})\big{)}^{T+1}\big{(}\log(n)+\log^{2.5}(\frac{1}{\epsilon})\big{)}\Big{)},$
@@ VII.1 Classical Algorithm and Quantum Advantage
$D_{i}(x)\cdot x=\sum_{\alpha=1}^{K}\sum_{k=1}^{p}\Big{(}\prod_{j=1,j\neq k}^{p}x^{T}A_{j}^{\alpha,i}x\Big{)}A_{k}^{\alpha,i}\cdot x,$
which takes a further complexity $\mathcal{O}(Kp^{2}ns)$. Finally, solving the linear equation (4) takes a further $\mathcal{O}(n^{3})$ time. Therefore, summing up the steps, the total time required for a total of $T$ iteration steps is:
$\mathcal{O}\Big{(}T(n^{3}+Kp^{2}ns+n^{p+1}s)\Big{)}.$ (34)
For large $p$, e.g., $p\geq 3$, the task is practically intractable on a classical computer if we manually follow the above steps. It is well-known that per Newton’s method, if the initial solution $x_{0}$ is sufficiently close to the real zero of given equations, then the solution converges quadratically fast, which means that the total number of iteration times $T\approx\mathcal{O}(\log\log(1/\epsilon))$ where $\epsilon$ is the desired error. The above running time indicates that the speedup of the quantum algorithm is superpolynomial, almost exponential, with respect to the number of variables.
We remark that the problem of solving a system of polynomial equations was, in fact, Smale’s 17th problem, where he asked if there exists a polynomial-time classical algorithm that outputs an approximate zero of such equations, given that these equations could be computed at any given point on demand. The answer turned out to be positive after multiple efforts, which are summarized and finally settled in the seminal work burgisser2011problem . The best-known running time, as mentioned in burgisser2011problem is $\mathcal{O}(n^{\log\log(n)})$, which is almost linear in the number of variables $n$. We emphasize an important subtlety, namely, they assume the ability to compute any given polynomial via certain means; however, the structure of these functions, as well as the complexity of how to compute them efficiently is not taken into account. Therefore, the time complexity of our quantum algorithm, as stated in Theorem 1, should be compared to the classical one shown in Eqn. (34).
Newton’s method typically works well given that we randomly initialize a vector within the vicinity of the real solution. Therefore, it is reasonable that Newton’s method (both classical and quantum versions) will succeed after a few iterations, e.g., when $T$ is small. In this case, a huge quantum speedup is obtained. Now, we move on to discuss a way to achieve better initialization from a quantum algorithmic perspective.
@@ VII.2 Quantum Approach for Initialization
Aside from the particular algebraic form that we impose, the behavior of functions of interest is quite critical. We recall that the Newton method is usually effective when the initially randomized solution is closed to the real one, as the method would succeed after after a few iterations. Therefore, it is quite important to sort out a way to prepare such a good initialization. A general strategy to prepare such a thing is not known to us. Here, we aim to propose a way to accelerate such a task, inspired by the way that each of the functions $f_{i}$ can be computed quickly using a quantum computer.
@@ VIII Motivating Examples And Potential Applications/Extensions
VIII.0.1 Nonlinear Optical Process/ Gross-Pitaevskii Equation
@@
where $\Delta t,\Delta x$ and $V_{j}^{n}$ are time step, grid spacing, and potential at the grid point $(x_{j},t_{n})$. The above equation at different grid points represents a coupled nonlinear algebraic equation, which could be solved using the classical Newton method to obtain a vector solution that contains all $\{\psi_{j}^{n}\}$ at different grid points. However, we point out two issues. First, our method is only applied to a nonlinear system where each equation is a homogeneous polynomial of even degree. Second, the above equation (Eqn. 43) actually contains both linear and nonlinear terms. In order to integrate our method, an adjustment is hence required. Apparently in Section VII.3 we showed how to deal with this scenario, but here we point out that it can be alternatively done in a simpler manner.
@@
$g=\begin{pmatrix}g_{1}=a_{1}x+b_{1}y+c_{1}z\\ g_{2}=a_{2}x+b_{2}y+c_{2}z\\ g_{3}=a_{3}x+b_{3}y+c_{3}z\end{pmatrix},$ (53)
@@
Now, we observe that the nonlinear equation in Ref. 43 contains both linear and nonlinear terms. Therefore, if we separate the function on the left-hand side of Eqn. (43) into linear and nonlinear parts, then the Jacobian of such equations, which basically contains gradients of corresponding functions, is the sum of the Jacobians from the constituting functions. The Jacobian of the linear constituent function is easily computed based on the above example. The Jacobian of the nonlinear constituent is computed using the algorithm outlined in our discussion above.
@@
VIII.0.2 Lotka-Volterra Equations for Modeling
@@
VIII.0.3 Intersection Of Algebraic Vaieties
Algebraic variety is a fundamental object in algebraic geometry (see Figure 2 for some illustrating examples). Formally, it is a basically a set of solutions to a system of polynomial equations over a field.
$x$$y$
(a) Affine space $\mathbb{A}^{2}$: basically four points on $x-y$ plane
$x$$y$
(b) Algebraic curve: a unit circle
$x$$y$
(c) Algebraic surface: an elliptic curve
Figure 2: Some example of algebraic varieties. Figure 2(b) features a one-dimensional algebraic variety. Figure 2(c) features two-dimensional algebraic variety.
In algebraic geometry, one of the main theme is studying geometric shapes defined by polynomial equations. Given multiple varieties, then pictorially, one can imagine that their intersection form exactly a system of nonlinear algebraic equation. Solving for such intersection for a high dimensional/ complicated variety is classically difficult, and in fact, as we pointed out in the work, for degree $\geq 3$, solving such equation is even intractable. At the same time, many important application tasks are based on algebraic geometry concepts, in particular, algebraic variety. In robotics and computer-aided design (CAD) mortenson1997geometric , shapes are represented using algebraic varieties, and their intersection could model collision. Thus, finding the intersection properly could potentially be used for collision avoidance strategy in robot motion planning. Furthermore, in the area of robot vision and image processing szeliski2022computer , algebraic varieties, such as algebraic curves and surfaces are also used for representing object and and scenes in images. Stemming from our efficient quantum algorithm, we envision a future research that designs quantum algorithm for object recognition, feature extraction, etc from given algebraic representation of objects and scenes, as mentioned.
@@ IX Conclusion
In this article, we have developed a quantum algorithm for solving nonlinear algebraic equations of arbitrary polynomial type. By employing well-known quantum tools, such as the unitary block encoding method, an iterative procedure building upon classical Newton’s method is constructed to drive the initially randomized guess to the close-to-real solution, which is expected to succeed after only a few iteration steps. Additionally, we have provided two examples from physical, biophysics and algebraic geometry contexts to suggest the potential application of our method, as well as suggesting another way to extend our method to deal with a more specific setting.
@@ X Acknowledgement
We thank Xianfeng Gu and Nengkun Yu for their useful discussions. This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under Award Number DE-SC-0012704. We also acknowledge the support by a Seed Grant from Stony Brook University’s Office of the Vice President for Research and by the Center for Distributed Quantum Processing.
@@ Appendix C Proof of Lemma 7
According to the definition 1, the unitary $U_{m}U_{p}$ is the block encoding of the matrix: ${x_{0}^{2p-1}J}/{(\sqrt{n}ps)}$. One can reasonably expect that if the value of $x_{0}$ is very small, then it will incur a substantial running time. A better way to overcome this issue is to change the state |$\ket{\bf 0}_{A}$ in the above construction to some states that are close to any temporal solution $x_{k}$ during the iteration. As we have mentioned, suppose the initially guessed solution $x_{0}$ is within the vicinity of the actual solution, then the overlaps between any pairs of $x_{0}$ and temporal solution $x_{k}$ (at $k$-th iteration step) are supposed to be sufficiently high, e.g., lower bounded. Therefore, in the above procedure, one simply needs to let unitary $U$ act on the register $A$ to transform the state |$\ket{\bf 0}_{A}$ to the state |$\ket{x_{0}}$ (or $x_{0}$ as we denoted before) and proceed accordingly. Then the factor $\gamma$ in Lemma 7 would be the overlap $|x_{0}^{T}x_{k}|$, which could be estimated using amplitude estimation [32, 33]. This step only incurs a constant cost overhead and, hence, doesn’t increase the overall running time.